# On the foundation of quantum decision theory

### v3

@@ 4 Concluding remarks
Quantum decision theory seems to be applicable in many cases where decisions are to be made, primarily but not only, in the cases that Kahneman (2011) calls decisions of type 1, fast decisions. When a prior is available, an alternative is a Bayesian decision theory. This theory will follow the Kolmogorov rules, and apart from the existence of a prior, the theory does not require any special assumptions.
In this article I have concentrated on decisions made by a single person. however, the theory can be extended to joint decisions made by a group of communicating persons. Then these persons must have common ideals, as modelled by a perfectly rational higher being $D$. The ideals must be common in the setting determined by the relevant decision variables.
@@
As a basis for quantum decision theory as discussed here, lies the idea of two complementary decision variables at the same time in the mind of some person or in the joint minds of a group of communicating persons. I repeat that two different variables are complementary if they are maximal as accessible variables. Another essential assumption is the existence of ideals that are perceived as perfectly rational, modelled by a hypothetical ‘ideal higher being’.
The quantum probabilities may be thought of as relative to this ‘ideal higher being’, hence they must be looked upon as absolute probabilities, not conditioned in any way. This is confirmed by empirical findings in the quantum decision literature.
There are cases where powerful political leaders may be roughly modelled as such ‘ideal higher beings’, relative to the influence they have on other people and their decisions. A fundamental requirement should then be that the leaders obey certain basic humanitarian conditions, including respect for other people’s lives. Unfortunately, this goal is not always satisfied in practice.
Examples where the above two assumptions hold, may be found within the mind of any person, but also between persons and between groups of persons.
The assumptions may in certain situations be satisfied by many different pairs of decision variables, leading to mental difficulties in the first case and to partly complementary world views in the last two cases. Then in these cases it might imply deep conflicts between those persons or groups if some necessary conditions are there. We can see many examples of this phenomenon in the world today.
5 Aknowledgments

### v4

@@ 5 Concluding remarks
Quantum decision theory seems to be applicable in many cases where decisions are to be made, primarily in the cases which Kahneman (2011) calls decisions made through fast thinking, thinking of type 1. When a prior is available, an alternative is a Bayesian decision theory. This theory will follow the Kolmogorov rules, and apart from the existence of a prior, the theory does not require any special assumptions.
In this article I have concentrated on decisions made by a single person. But the theory can be extended to joint decisions made by a group of communicating persons. Then these persons must have common ideals, as modelled by a perfectly rational higher being $D$. The ideals must be common in the setting determined by the relevant decision variables.
@@
As a basis for quantum decision theory as discussed here, lies the idea of two complementary decision variables at the same time in the mind of some person or in the joint minds of a group of communicating persons. I repeat that two variables are here called complementary if they are really different (not in one-to-one correspondence) and maximal as accessible variables. Another essential assumption is the existence of ideals that are perceived as perfectly rational, modelled by a hypothetical ‘ideal higher being’.
The quantum probabilities may be though of as relative to this ‘ideal higher being’, hence they must be looked upon as absolute probabilities, not conditioned in any other way than by our past knowledge. This is confirmed by empirical findings in the quantum decision literature.
There are cases where powerful political leaders may be roughly modelled as such ‘ideal higher beings’, relative to the influence they have on other people and their decisions. A fundamental requirement should then be that the leaders obey certain basic humanitarian conditions, including respect for other people’s life. Unfortunately, this goal is not always satisfied in practice.
Examples where the assumptions behind decisions lead to conflicts, may be found within the mind of any person, but also between persons and between groups of persons.
These assumptions may in certain situations be satisfied by very many different pairs of decision variables, perhaps leading to mental difficulties in the first case and to partly complementary world views in the last two cases. Then in these cases it might imply deep open conflicts between those persons or groups if some necessary conditions are there. We can see many examples of this phenomenon in the world today.
It is a hope that ideas related to this article may lead to an increased focus on making the best decisions for all people, for political leaders and in a global perspective, the best decisions in order to obtain peace on this earth and a viable future for all human beings.
Aknowledgements
@@
Conflicts of interest statement
There are no conflicts of interest.
Data availability statement
There are no data in this article.